OpenStudy (anonymous):

Jim can fill a pool carrying buckets of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 ½ hours. How quickly can all three fill the pool together?

OpenStudy (anonymous):

Jim can fill a pool carrying buckets of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 ½ hours. How quickly can all three fill the pool together?

OpenStudy (anonymous):

please help me to solve this one

OpenStudy (saifoo.khan):

trying..

OpenStudy (anonymous):

The trick to these is to reverse the way you look at them. Tony can fill 2 pools in 3 hours. Sue can fill 4 pools in 3 hours, and Jim can fill 6 pools in 3 hours. (I used 3 hours so that they all work out to whole numbers) So, combined, they can fille 12 pools in 3 hours...or 4 pools per hour. This is the same a 1 pool in 15 minutes.

OpenStudy (amistre64):

J does 1/30 in 1 min S does 1/45 in 1 min T does 1/90 in 1 min add them all up

OpenStudy (amistre64):

3/90 2/90 1/90 ----- 6/90 ; flip it to get 90/6 minutes

OpenStudy (anonymous):

i don't understand dear

OpenStudy (anonymous):

mine, or amistre's?

OpenStudy (amistre64):

lets assume it takes you 5 minutes to walk 1 mile; how many miles do you walk in 1 minute?

OpenStudy (anonymous):

12

OpenStudy (amistre64):

1/5 of a mile in 1 minute, "plus" 1/5 of a mile in 1 minute, "plus" 1/5 of a mile in 1 minute, "plus" 1/5 of a mile in 1 minute, "plus" 1/5 of a mile in 1 minute, "plus" --------------------------- 5/5 of a mile in 5 minutes .... right?

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

in other words; you figure out how much work is down in "one" interval if it takes J 30 minutes to do the whole job; how much is done in 1 minute?

OpenStudy (anonymous):

1/30

OpenStudy (amistre64):

good; now lets do the same for S, it takes her 45 minutes to do the job, how much is done in 1 minute by her?

OpenStudy (anonymous):

1/45

OpenStudy (amistre64):

excellent; now if we add their contributions together, we get the total amount of work done in 1 minute if they work together right?

OpenStudy (anonymous):

so what is equation for this ?

OpenStudy (anonymous):

1/30+1/45+1+90=?

OpenStudy (amistre64):

1/J + 1/S + 1/T = total amount of work done by J+S+T in 1 minute right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

since and hour and a half equals 90 minutes, we get them all working in minutes ..

OpenStudy (amistre64):

1/30 = J 1/45 = S 1/90 = T -------- 6/90 total ; remember how to add fractions

OpenStudy (amistre64):

since the whole pool divided into 90 parts and 6 parts are filled each minute if they work together... 6*t = 90 t = 90/6